3. What does Somers d value of -0.70 specifically tell you about the nature of the relationship between the two variables in a crosstab table?

4. Calculate the degrees of freedom associated with a 5 by 4 crosstab table. Show the solution.
Answer #1

3)

Somers d = -0.7 means that there was an inverse relationship between two variables.

Somers’d is a non-parametric measure of the strength and direction of association that exists between an ordinal dependent variable and an ordinal independent variable.

4) Given that

Number of rows = 5

Number of columns = 4

Degrees of freedom = (rows-1)(columns-1) = (5-1)(4-1) = 12

Degrees of freedom = 12